Suppose a power line produces a 5.1×10−4-T peak magnetic field 60 times each second at the location of a neuron brain cell of radius 6.0 ×10−6 m.

Determine the maximum magnitude of the induced emf around the perimeter of this cell during one-half cycle of magnetic field change.
Answer #1

Flux = B.A

E = d(flux)/dt

E = dB/dt * A

E = 5.1*10-4 * 60 * \pi * r2

E = 5.1*10-4 * 60 * \pi * (6*10-6)2

E = 3.46 * 10-12 V​